OpenStudy (anonymous):

the line through (1,7) and (-6,7) in point-slope form?

OpenStudy (agreene):

point slope form is: \[y-y_1=m(x-x_1)\] with \[m=\frac{\Delta y}{\Delta x}\implies \frac{y_2-y_1}{x_2-x_1}\] you are given two points in the terms: \[(x_1,y_1);(x_2,y_2)\] plug your points into the m= equation to solve for it, then plug your points and slope in to the general form.

OpenStudy (anonymous):

but the slope for those coordinates is 0

OpenStudy (anonymous):

so the answer would be y-7=0(x-1) ?